# Experimental Study of Underwater Acoustic Reconfigurable Intelligent Surfaces with In-Phase and Quadrature Modulation

### v1

@@ VII Conclusion and Future Work
In this paper, we explored the design of UA-RIS to enhance the range and speed of underwater acoustic communications. We addressed several challenging issues critical to system implementation, including wide dynamic range and low frequency of acoustic signals. To optimize performance, the proposed UA-RIS strategically aligns reflection units based on the angles of incident waves and their intended directions, utilizing IQ modulation to shape the reflected waves. This approach can precise control over the phase and amplitude of the signals, enabling advanced beam-steering capabilities.
@@
Results from tank experiments demonstrated the capability of acoustic reflectors to finely tune the phase and amplitude of signals by adjusting the load impedance. Additionally, we evaluated a UA-RIS setup comprising 24 reflection units during lake trials, which proved effective in significantly enhancing or diminishing signal strength in designated directions. This adjustment improved receiving SNR and minimized interference at specific locations. Experimental results suggest that, by deploying the UA-RIS close to the sender, the communication range in deep and shallow water environments can be extended by 28% and 46%, respectively. When the distance between communicating parties remains constant, data transmission rates employing modulation schemes like FSK or PSK can be increased by an average of 63.8%, and in some cases, up to 96% with the support of UA-RIS. In addition, moiving the UA-RIS near the receiver extends the communication range in both deep and shallow water environments by 40.6% and 66%, respectively. Furthermore, with the distance fixed, situating the UA-RIS near the receiver not only improves data rates by an average of 80.3% but can also achieve increases of up to 163% under specific conditions.
@@
UA-RIS technology is still in its early stages, its real-world deployment requires further development. In particular, integrating this new technology into existing systems necessitates critical modifications to current medium access control protocols. Adapting these protocols will enable the seamless integration of UA-RIS into established underwater network infrastructures, a topic we intend to investigate in our future research.

### v2

@@ VI Conclusion and Future Work
In this paper, we explored the design of UA-RIS to enhance the range and speed of underwater acoustic communications. We addressed several challenging issues critical to system implementation, including wide dynamic range and low frequency of acoustic signals. To optimize performance, the proposed UA-RIS strategically aligns reflection units based on the angles of incident waves and their intended directions, utilizing synthetic reflection method to shape the reflected waves. This approach can precise control over the phase and amplitude of the signals, enabling advanced beam-steering capabilities.
@@
Results from COMSOL Multiphysics simulation and tank experiments demonstrated the capability of acoustic reflectors to finely tune the phase and amplitude of signals by adjusting the load impedance. Additionally, we evaluated a UA-RIS setup comprising 24 reflection units during lake trials, which proved effective in significantly enhancing or diminishing signal strength in designated directions. This adjustment improved receiving SNR and minimized interference at specific locations. Experimental results using a CW signal as the source indicate that deploying the UA-RIS near the transmitter can extend the communication range by up to 28% and 46% in deep water and shallow water environments, respectively. When the distance between communicating parties remains constant, data transmission rates employing modulation schemes like FSK or PSK can be increased by an average of 63.8%, and in some cases, up to 96% with the support of UA-RIS. In addition, moving the UA-RIS near the receiver extends the communication range in both deep and shallow water environments by 40.6% and 66%, respectively. Furthermore, with the distance fixed, situating the UA-RIS near the receiver not only improves data rates by an average of 80.3% but can also achieve increases of up to 163% under specific conditions.
@@
UA-RIS technology is still in its early stages, its real-world deployment requires further development. In particular, integrating this new technology into existing systems necessitates critical modifications to current medium access control protocols. Adapting these protocols with different modulation schemes (e.g., FSK, PSK, and OFDM) will enable the seamless integration of UA-RIS into established underwater network infrastructures, a topic we intend to investigate in our future research.